Two forces of 60 N east and 30 N west act on a 300kg mass. if friction is neglected, what is the resulting acceleration of the mass?
1
Expert's answer
2019-05-06T10:08:58-0400

Net force is


"\bold{F=F_1+F_2}"


"F=F_1-F_2=60-30=30\ N"

directed eastwards.

The resulting acceleration of the mass:


"a=\frac{F}{m}=\frac{30}{300}=0.1 \frac{m}{s^2}"

directed eastwards.